# Superconducting spintronics with electron symmetry filtering and interfacial spin-orbit coupling

### v1

@@ III.4 TMR enhancement with applied bias
As we have already mentioned above, the orbital symmetry-controlled TMR was first predicted and then experimentally confirmed about two decades agoButler2001; Mathon2001. Today, it is probably the most notable phenomenon in the whole spintronics field, finding applications from magnetoresistive random access memories, logic gatesNey2003 and magnetic sensors Lenz2006. TMR was first prediced to reach values exceeding $1000\%$ at $V=0$ Butler2001; Mathon2001. Unfortunately, this is not possible in experimental setups due to effects such as diffusion or roughness at atomic interfaces or emerging electronic states near the surfaces, which only allow for a $\sim 200-600\%$ TMR in realistic casesIkeda2008.
@@
A big challenge that TMR has to overcome in MTJs is its degradation with applied bias Parkin2004; Yuasa2004; Waldron2006; Guerrero2007; Herranz2010. In non-epitaxial devices where the electron tunneling is incoherent, this is due to electrons with extra energy above $E_{F}$, (so-called “hot electrons” Parkin1997), that are usually produced when bias or temperature are increased. These hot electrons then produce magnon excitations near the interface where the tunneling takes place, which in turn results in spin-flip events during tunneling Moodera1995. The parallel configuration of the F electrodes (in standard F/F junctions) is not greatly affected by this since the spin-flips do not result in extra electrons tunneling, but the same effect has a much bigger influence on the bias dependent conductance in the anti-parallel configuration, as here spin-flips allow the electrons to pass through the barrier, so the TMR (which measures the relation of the P and AP configuration conductances) is also reduced. In epitaxial junctions, coherent transmission between electronic bands is able to conserve the orbital symmetries in the tunneling processGuerrero2007. When a bias is applied to the MTJ, this band-to-band transport has a stronger effect for the anti-parallel configuration than for the parallel one, again reducing the TMRWaldron2006.
Interestingly. the conductance and TMR of N/F/F (V/MgO/Fe/MgO/FeCo) junctions shows unusual behavior under applied biasGonzalezRuano2021-HB-TMR: their TMR steadily increases from below 50% to above 100% with an applied bias below 0.5 V in a broad temperature range, from cryogenic to room temperature (Figure 6d). We have found that SOC was the main factor behind this effect, as it controls the tunneling of the N/F (V/MgO/Fe) part of the junctions while the F/F part is more standard, depending on the magnetic orientation of the two layers. In Figure 6d, we show a comparison of the TMR vs bias in a F/F (a) and a N/F/F (c) junction. As expected, TMR is higher at zero bias and reduces with applied bias in the F/F junction, while it is enhanced with applied bias in the N/F/F one, at least up to $V=0.5$ V.
@@
$\Delta V=\left|V\right|\times\frac{R_{AP}-R_{P}}{R_{P}}$ (1)
@@
is also an important factor. In our junctions, it improves the values of F/F junctions with similar structure at biases above $V=0.5$ V, as we demonstrate in Figure 6e. It is worth noting that, to the best of our knowledge, its value at room temperature (above $\Delta V=0.8$ V) sets a record in spintronicsTiusan2006. At low applied biases it is still larger in F/F junctions, but in these junctions it rapidly decreases with applied bias, contrasting with the steady increasing in N/F/F ones.
@@
To find the explanation for this unusual behavior, we can first look at Figure 6b and notice a strong conductance increase that takes place with applied bias in the N/F junctions. Due to the aforementioned symmetry filtering, these MTJs are rather resistive at low biases. To give a comparison without the MgO-induced conductance bottleneck, control samples with an Au layer instead of a Fe one have a thousand times larger conductance when the lateral size is the same (see Ref.GonzalezRuano2025, Supplemental Material). For the N/F/F junctions we show the conductance in both the parallel and anti-parallel configurations of the F/F part in Figure 6c. Here one can observe how there is a range of biases where both conductances diverge. In Figure 6g, we sketched a region of surface states in the vanadium layer, which as we mentioned present electronic symmetries different than the rest of the layer.
@@
These states were found with ab-initio calculations similar to the ones described in Section III.2, and are crucial for understanding the TMR vs bias behavior, as these states allow us to explain the high zero bias resistivity of V/MgO/Fe: in vanadium, there is only one $\Delta_{2}$ electronic band crossing the Fermi level in the normal incidence direction. However, this symmetry does not exist at the Fermi level on the other side of the barrier, in the Fe electrode (and if was, the MgO barrier would stop the transmissionButler2001). The low bias transport takes place only when $\Delta_{2}$ electrons change their orbital symmetry to $\Delta_{1}$ in the region with these surface states. A similar phenomenon was already published, where Fe/V/Fe/MgO/Fe junctions had much higher TMR than similar ones without the V layerFeng2009.
@@
The DFT calculations are similar to the ones explained in the previous section. This time, instead of probing the possible presence of SOC in our system, we wanted to look for mechanisms that could allow the observed transport features to take place. What the new calculations revealed was a peak in the vanadium surface density of states near the MgO, just below $V=100$ mV, which we found to agree with previous scanning tunneling spectroscopy experimentsBischoff2001. The crucial part is that the calculations showed that these surface DOS would be enhanced if we added interfacial SOC. The reason for this is that, without SOC, the Fermi level falls right inside a gap for $\Delta_{1}$ states, which is no longer present in with SOC due to scattering processes allowing $\Delta_{1}$ and $\Delta_{2}$ to mix, as sketched in Figure 5a.
With these calculation results, we are able to model the system with a relatively simple approximation that satisfactorily reproduced the TMR vs bias experimental results. The starting point for this model is to notice that the change in magnetic configuration from P to AP not only affected the total resistance, but also modified the share of voltage drop that was taking place around each of the two MgO barriers of the N/F/F junctions. Taking into account the shape of the experimentally measured conductance curves and the new DOS peak explaining their shape, we were able to accurately reproduce the TMR($V$) of the junction from what we considered to be two nonlinear, in-series components: the V/MgO/Fe and Fe/MgO/Fe/Co parts, for which we had several separate conductance vs. $V$ curves (or rather, the $I(V)$ ones) from experiments in different types of MTJs. It is worth noting that a crucial step here that makes this possible is that the Fe layer where we make the separation of the two “sub-junctions” has a thickness of 10 nm, which prevents coherent tunneling through both barriers as it is larger than the coherence length of the electrons for both spin populations, forcing them to undergo a sequential tunnelling through each of the barriers separately Tao2019.
@@
IV SOC-induced effects in S/F systems: low temperature experiments
This chapter explains how the generation of spin-triplet Cooper pairs in epitaxial S/SOC/F systems with spin and orbital symmetry filtering has been experimentally probed. Some effects discussed below have been anticipated theoretically. These predictions include (i) magnetoanisotropic Andreev reflection Hogl2015 and (ii) superconductivity-induced changes in MCAJohnsen2019. As further indirect evidence, we also use above-gap conductance anomalies (MacMillan resonances MRR1) to verify the possible contributions of SOC and magnetic spin textures to the generation of LRTsVisani2012.
IV.1 Magneto-anisotropic Andreev reflection
In this section, we will see how low bias magnetoresitance in all-epitaxial S/F MTJ reveals a substantial modification below the superconducting transition, gaining a three orders of magnitude higher MR anisotropy. The non-volatile magnetic configurations (IP and OOP) allow us to rule out orbital and vortex effects and identify the SOC origin of the observed MR. Such MR reaches $\sim 20\%$ without an applied magnetic field, and is further increased for large magnetic fields Martinez2020 .
In MTJs with SOC, one only needs one ferromagnetic layer to achieve MR. This is the so-called tunneling anisotropic MR (TAMR) Fabian2007. Unfortunately, its usual values are of around $1\%$ Lu2012, which for most practical purposes is useless. In the S/F MTJs under study, we found a huge increase in the MR of the Fe layer when V cools down below $T_{C}$, becoming superconducting. As we will see, interfacial SOC plays a central role again, suggesting the presence of magnetization-dependent LRTs formation. Notably, the materials that these MTJs are composed of are readily available and well-tested in commercial uses, since they are based on Fe/MgO junctions Parkin2004; Yuasa2004; Cascales2012.
To see these effects, first one has to search for a distinct feature of interfacial SOC: above $T_{C}$, the influence of interfacial SOC has observable effects on the TAMR of N/F junctions, resulting in magnetization-dependent conductance Fabian2007. This effect by itself is too faint for any potential uses (usually $\text{TAMR}=[G(0)-G(\pi/2)]/G(\pi/2)\approx 0.01\%$)Martinez2020. However, exploring the conductance in the same junction under $T_{C}$ reveals a much higher anisotropy. In the superconducting regime, the counterpart for TAMR is the magneto-anisotropic Andreev reflection (MAAR) Hogl2015, which has the same formula, where the conductance is instead taken to be within the superconducting gap:
$\text{MAAR}=\dfrac{G(0)-G(\theta)}{G(\theta)},$ (2)
Appart from the MAAR values, we can extract more information from the experimental data Martinez2020. Perhaps one of the most important parts is the transparency of the tunnel barrier. To model the experimental data, we used a BTK model Hogl2015; Blonder1982 where the transparency of the barrier in the normal state is given by $1/(1+Z^{2})$ (here, $Z=0$ means perfect transparency, while $Z\gg 1$ is the tunnel junction limit). Therefore, the stronger (less transparent) the barrier is, the more the conductance will decrease Blonder1982. In this regard, Figure 7a shows a comparison of (normalized) conductances of the same sample below and above $T_{C}$ ($T=0.3$ K and $T=10$ K). From the curves, the conductance change can be used to extract a transparency which turns out to fall in an intermediate regime (RefMartinez2020 discusses this in more detail), in between typical values among those of point contacts Soulen1998; Nadgorny2012 and spin-polarized tunneling Zutic2004; Parker2002; Parkin2004.
@@ IV.1 Magneto-anisotropic Andreev reflection
Andreev reflections are also highly susceptible to SOC, in particular for high spin polarizations Zutic2004 such as in our system. This is because, in the limit of full spin polarization, conventional Andreev reflection is not allowed, so the conductance drops to zero. This totally changes when interfacial SOC appears, as it allows for an analogue process called spin-flip Andreev reflection. In the case of V/MgO/Fe junctions, the Fe/MgO barrier has a very high effective spin polarization thanks to the symmetry filtering Zutic2004; Parkin2004; Yuasa2004. On top of this, it also has a structural inversion symmetry breaking at the interface (as the atomic lattice changes its periodicity and potential), which is the main underlying cause of the Rashba SOC Zutic2004; Fabian2007. In Figure 7a, we show how the conductance is only mildly suppressed if we compare it to what one could expect for a similarly spin-polarized interface ($\sim 75\%$), which suggest that the spin-flip Andreev reflection process explained above is responsible for the relatively high values of $G(V=0)$ observed.
[…]
To complete this picture, Figure 7e,f depict the zero bias conductance during an IP (0°) to OOP (90°) magnetization rotation, where the phenomenological model that was used to understand the results is also pictured with solid lines. The reasonable fit allowed us to establish with more certainty Rashba SOC as the main contributor to the MAARMartinez2020.
@@ IV.2 Superconductivity-induced change in magnetic anisotropy
Due to the characteristic energies of the magnetic and superconducting order parameters, it’s common to see from experimental results that the magnetism affects (or totally disrupts) superconductivity in a neighbouring material. The opposite is not common, since the magnetic exchange field in ferromagnets is more energetic than superconducting correlations ($\sim 10^{3}$ K vs $\sim 10^{1}$ K respectively). However, an important magnetic property of epitaxial materials, the magneto-crystalline anisotropy (MCA) does have a energy scale which is similar to the superconducting gap, and therefore it’s more susceptible for a mutual interaction with superconductivity. This is precisely what we will show in this section, where the results let us observe how superconductity changes the easy magnetization axis of the adjacent ferromagnetic layer.
This phenomenon was known from theoretical studies since 2019 Johnsen2019, but was not experimentally verified until a couple years later gonzalez-prb-20; gonzalez-scirep-21. The underlying physics involve, once again, Rashba SOC and LRT generation. In this case, in particular, the crucial part is that the generation rate of LRTs is sensitive to the relative direction of the magnetization and the SOC field. When the superconducting transition takes place, and to minimize the total energy in the whole system (V/MgO/Fe), the magnetization re-orients itself to a direction where less LRTs are formed (as they are energetically expensive), even if that means aligning with the hard magnetization axis. Superconductivity acts as a “switch” for the magnetization, which could be a really useful approach designing future devices where magnetization control is still a significant challenge to address Banerjee2014; Baek2014; Gingrich2016; Satchell2021.
For clarity, this section is separated in two parts, both studying the S/F/F junctions that were described in Figure 1c. First, we will verify the superconductivity induced modification of the IP magnetic anisotropy, and after that we will discuss modification of the OOP anisotropy.

### v2

@@ III.4 TMR enhancement with applied bias
As we have already mentioned above, the orbital symmetry-controlled TMR was first predicted and then experimentally confirmed about two decades agoButler2001; Mathon2001. Today, it is probably the most notable phenomenon in the whole spintronics field, finding applications from magnetoresistive random access memories, logic gatesNey2003 and magnetic sensors Lenz2006. TMR was first prediced to reach values exceeding $1000\%$ at $V=0$ Butler2001; Mathon2001. Unfortunately, this is not possible in experimental setups due to structural defects and other effects, such as diffusion or roughness at atomic interfaces or emerging electronic states near the surfaces, which only allow for a $\sim 200-600\%$ TMR in realistic casesIkeda2008.
@@
Another major challenge that TMR has to overcome in MTJs is its degradation with applied bias Parkin2004; Yuasa2004; Waldron2006; Guerrero2007; Herranz2010. In non-epitaxial devices where the electron tunneling is incoherent, this is due to electrons with extra energy above $E_{F}$, (so-called “hot electrons” Parkin1997), that are usually produced when bias or temperature are increased. Then, these hot electrons produce magnon excitations near the interface where the tunneling takes place, which in turn result in spin-flip events during tunneling Moodera1995. The parallel configuration of the F electrodes (in standard F/F junctions) is not greatly affected by this since the spin-flips do not result in extra electrons tunneling, but the same effect has a much bigger influence on the bias-dependent conductance in the AP configuration, as here spin-flips allow the electrons to pass through the barrier, so the TMR (which measures the relation of the P and AP configuration conductances) is also reduced. In epitaxial junctions, the coherent transmission between electronic bands is able to conserve the orbital symmetries in the tunneling processGuerrero2007. When a bias is applied to the MTJ, this band-to-band transport has a stronger effect for the AP configuration than for the P one, reducing the TMRWaldron2006.
Interestingly, the TMR of N/F/F (V/MgO/Fe/MgO/FeCo) junctions shows an unusual behavior under applied biasGonzalezRuano2021-HB-TMR: it steadily increases from below 50% to above 100% with an applied bias below 0.5 V in a broad temperature range, from cryogenic to room temperature (Figure 6d). We have found that SOC was the main factor behind this effect, as it controls the tunneling of the N/F (V/MgO/Fe) part of the junctions (this same mechanism will still take place, accompanied by other processes, when we discuss transport in the superconducting regime in the following sections) while the F/F part is more standard, depending on the magnetic orientation of the two layers. In Figure 6d, we show a comparison of the TMR vs bias in a F/F (a) and a N/F/F (c) junction. As expected, TMR is higher at zero bias and reduces with applied bias in the F/F junction, while it is enhanced with applied bias in the N/F/F one, at least up to $V=0.5$ V.
@@
$\Delta V=\left|V\right|\times\frac{R_{AP}-R_{P}}{R_{P}},$ (2)
@@
is also an important factor. In the equation above, $R_{P}$ ($R_{AP}$) is the value of the electrical resistance in the P (AP) magnetic configurations. In our junctions, it improves the values of F/F junctions with similar structure at biases above $V=0.5$ V, as we demonstrate in Figure 6e. It is worth noting that, to the best of our knowledge, its value at room temperature (above $\Delta V=0.8$ V) sets a record in spintronicsTiusan2006. At low applied biases, $\Delta V$ is even larger in F/F junctions, but it rapidly decreases with applied bias, in contrast to the steady increase in N/F/F junctions.
@@
To find the explanation for this unusual behavior, we can first look at Figure 6b and notice a strong conductance increase that takes place with applied bias in the N/F junctions. Due to the aforementioned symmetry filtering, these MTJs are rather resistive at low biases. To give a comparison without the MgO-induced conductance bottleneck, control samples with an Au layer instead of a Fe one have a thousand times larger conductance when the lateral size is the same (see Ref.GonzalezRuano2025, Supplemental Material). For the N/F/F junctions, we show $G$ in both the P and AP configurations of the F/F part in Figure 6c. Here one can observe how there is a range of biases where both conductances diverge. In Figure 6g, we sketched a region of surface states in the vanadium layer, which as we mentioned present electronic symmetries different than the rest of the layer.
@@
These states were found with ab-initio calculations similar to the ones described in Section III.2, and are crucial for understanding the TMR vs bias behavior, as these states allow us to explain the high zero bias resistivity of V/MgO/Fe: in vanadium, there is only one $\Delta_{2}$ electronic band crossing the Fermi level in the normal incidence direction. However, this symmetry does not exist at the Fermi level on the other side of the barrier, in the Fe electrode (if it would exist, the MgO barrier would stop the transmissionButler2001). The low bias transport takes place only when $\Delta_{2}$ electrons change their orbital symmetry to $\Delta_{1}$ in the region with these surface states. A similar phenomenon was already published, where Fe/V/Fe/MgO/Fe junctions had much higher TMR than similar ones without the V layerFeng2009.
@@
The density functional theory (DFT) calculations are similar to the ones explained in the previous section. What the new calculations revealed was a peak in the vanadium surface density of states near the MgO, just below $V=100$ mV, which we found to agree with previous scanning tunneling spectroscopy experimentsBischoff2001. The crucial part is that the calculations showed that these surface DOS would be enhanced if we added interfacial SOC. The reason for this is that, without SOC, the Fermi level falls right inside a gap for $\Delta_{1}$ states, which is no longer present in with SOC due to scattering processes allowing $\Delta_{1}$ and $\Delta_{2}$ to mix, as sketched in Figure 5a.
With these calculation results, we are able to model the system with a relatively simple approximation that satisfactorily reproduced the TMR vs bias experimental results. The final result is depicted in Figure 6f, where we can observe a nice fit to the experimental TMR($V$) curve of a N/F/F junction. Specifically, we solve GonzalezRuano2021-HB-TMR the nonlinear circuit equations by using Kirchhoff voltage and current laws (KVL, KCL): The KCL or charge continuity condition, $I_{1}=I_{2}=I$, allows us to determine the voltage drops $V_{1}$ and $V_{2(P/AP)}$ at each barrier from the individual $i(V)$ of each of them (measured experimentally). The KVL provides the total voltage drop on the serial device, $V=V_{1}+V_{2(P/AP)}$. The total conductance will be $G_{P/AP}(V)=I/V$, and then the corresponding TMR($V$) from Eq.1 can be calculated. We also used a parametrized $G_{1}(V)$ curve for the N/F part, adjusted to obtain the best fit of the final TMR($V$) curve for a real N/F/F junction. The result, shown in Figure 6f , accurately matches the TMR vs. bias in the N/F/F MTJs.
IV SOC-induced effects in S/F systems: low temperature experiments
This section explains how the generation of spin-triplet Cooper pairs in epitaxial S/SOC/F systems with spin and orbital symmetry filtering has been experimentally probed. Some effects discussed below have been anticipated theoretically. These predictions include: (i) magnetoanisotropic Andreev reflection Hogl2015 and (ii) superconductivity-induced changes in MCAJohnsen2019. As further indirect evidence, we also use above-gap conductance anomalies (MacMillan resonances MRR1) to verify the possible contributions of SOC and magnetic spin textures to the generation of LRTsVisani2012.
IV.1 Magneto-anisotropic Andreev reflection
In this section, we will see how the low bias MR in all-epitaxial S/F MTJs reveals a substantial modification below the superconducting transition, gaining a three orders of magnitude higher MR anisotropy. The non-volatile magnetic configurations (IP and OOP) allow us to rule out orbital and vortex effects and identify the SOC origin of the observed MR. Such MR reaches $\sim 20\%$ without an applied magnetic field, and is further increased for large magnetic fields Martinez2020 .
In MTJs with SOC, one only needs one ferromagnetic layer to achieve MR. This is the so-called tunneling anisotropic MR (TAMR) Fabian2007. Unfortunately, its usual values are of around $1\%$ Lu2012, which for most practical purposes is useless. In the S/F MTJs under study, we found a huge increase in the MR of the Fe layer when the V is cooled down below $T_{C}$, entering the superconducting state. As we will see, interfacial SOC plays a central role, suggesting the presence of magnetization-dependent LRTs formation. Notably, the materials that these MTJs are composed of are readily available and well-tested in commercial uses, since they are based on Fe/MgO junctions Parkin2004; Yuasa2004; Cascales2012.
To observe these effects, one must first identify a distinct feature of interfacial SOC: above $T_{C}$, the influence of interfacial SOC has observable effects on the TAMR of N/F junctions, resulting in magnetization-dependent conductance Fabian2007. This effect by itself is too faint for any potential uses (usually $\text{TAMR}=[G(0)-G(\pi/2)]/G(\pi/2)\approx 0.01\%$)Martinez2020. However, exploring the conductance in the same junction under $T_{C}$ reveals a much higher anisotropy. In the superconducting regime, the counterpart for TAMR is MAAR Hogl2015. Both have the same definition, but for MAAR the conductance is considered for biases within the superconducting gap:
$\text{MAAR}=\dfrac{G(0)-G(\theta)}{G(\theta)},$ (3)
Apart from the MAAR values, we can extract more information from the experimental data Martinez2020. Perhaps one of the most important parts is the transparency of the tunnel barrier. To model the experimental data, we used a BTK model Hogl2015; Blonder1982 where the transparency of the barrier in the normal state is given by $1/(1+Z^{2})$ (here, $Z=0$ means perfect transparency, while $Z\gg 1$ is the tunnel junction limit). Therefore, the stronger (less transparent) the barrier is, the more the conductance will decrease Blonder1982. In this regard, Figure 7a shows a comparison of (normalized) conductances of the same sample below and above $T_{C}$ ($T=0.3$ K and $T=10$ K). From the curves, the conductance change can be used to extract a transparency which turns out to fall in an intermediate regime (RefMartinez2020 discusses this in more detail), in between typical values among those of point contacts Soulen1998; Nadgorny2012 and spin-polarized tunneling Zutic2004; Parker2002; Parkin2004.
@@ IV.1 Magneto-anisotropic Andreev reflection
Andreev reflections are also highly susceptible to SOC, in particular for high spin polarizations Zutic2004 such as in our system. This is because, in the limit of full spin polarization, conventional Andreev reflection is not allowed, so the conductance drops to zero. This totally changes when interfacial SOC appears, as it allows for an analogue process called spin-flip Andreev reflection uti1999. In the case of the V/MgO/Fe junctions, the Fe/MgO barrier has a very high effective spin polarization thanks to the symmetry filtering Zutic2004; Parkin2004; Yuasa2004. On top of this, it also has a structural inversion symmetry breaking at the interface (as the atomic lattice changes its periodicity and potential), which is the main underlying cause of the Rashba SOC Zutic2004; Fabian2007. In Figure 7a, we show how the conductance is only mildly suppressed if we compare it to what one could expect for a similarly spin-polarized interface ($\sim 75\%$), which suggest that the spin-flip Andreev reflection process explained above is responsible for the relatively high values of $G(V=0)$ observed.
@@
To complete this picture, Figure 7e,f depict the zero bias conductance during an IP (0°) to OOP (90°) magnetization rotation, where the phenomenological model that was used to understand the results is also shown with solid lines. The reasonable fit allowed us to establish with greater certainty Rashba SOC as the main contributor to the MAARMartinez2020.
@@ IV.2 Superconductivity-induced change in magnetic anisotropy
Due to the characteristic energies of the magnetic and superconducting order parameters, it is common to see from experimental results that the magnetism affects (or totally disrupts) superconductivity in a neighbouring material. The opposite is not common, since the magnetic exchange field in ferromagnets is more energetic than superconducting correlations ($\sim 10^{3}$ K vs $\sim 10^{1}$ K respectively). However, an important magnetic property of epitaxial materials, MCA does have a energy scale which is similar to the superconducting gap, and therefore it is more susceptible for a mutual interaction with superconductivity. This is precisely what we will show in this section, where the results let us observe how superconductity changes the easy magnetization axis of the adjacent ferromagnetic layer.
This phenomenon was known from theoretical studies since 2019 Johnsen2019, but was not experimentally verified until a couple years later gonzalez-prb-20; gonzalez-scirep-21. The underlying physics involve Rashba SOC and LRT generation. In this case, in particular, the crucial part is that the generation rate of LRTs is sensitive to the relative direction of the magnetization and the SOC field. When the superconducting transition takes place, and to minimize the total energy in the whole system (V/MgO/Fe), the magnetization re-orients itself to a direction where less LRTs are formed (as they are energetically expensive), even if that means aligning with the hard magnetization axis. Superconductivity acts as a “switch” for the magnetization, which could be a really useful approach for designing future devices where magnetization control is still a significant challenge to address Banerjee2014; Baek2014; Gingrich2016; Satchell2021.
For clarity, this section is separated in two parts, both studying the S/F/F junctions described in Figure 1c. First, we will verify the superconductivity induced modification of the IP magnetic anisotropy, and after that we will discuss modification of the OOP anisotropy.